X^(2/3) - 3X(1/3) - 4 = 0
Answer by Alan3354(69443)

X^(2/3) - 3X(1/3) - 4 = 0
---------------------------
That's a quadratic in x^(1/3)
(x^(1/3) - 4)*(x^(1/3) + 1) = 0
x^(1/3) = 4, -1
x = 64
x = -1